Prove that the equation of the chord joining the points P(ct, c/t) and Q(cT, c/T) on
the rectangular hyperbola xy = c
2
is x + tT y = c(t + T). M is the midpoint of P Q
and P Q meets the x-axis at N. Prove that OM = MN, where O is the origin
1
Expert's answer
2020-03-13T10:19:28-0400

P (ct,c/t) and Q (cT,c/T)

Slope of line joining PQ is "\dfrac{(c/T – c/t)}{(cT – ct)}= – 1/Tt"

 Equation of the line using slope-point form is "(y – ct) = (– 1/cTt)(x – c/t)\\"

Simplifying we get "x + tTy = c(t + T)."


M is the midpoint of PQ. By midpoint formula,

"M \equiv (c(t + T)/2 , c(1/t + 1/T)/2)"

i.e."M(c(t + T)\/2 , c(t + T)\/2tT)"


By distance formula,

"OM = \sqrt{({\dfrac{{c(t + T)}}2 – 0})^2 + ({\dfrac{c(t + T)}{2tT} – 0})^2}"

Simplify to get OM = "[c(t + T)\sqrt{(1 + T^2t^2)] / 2}] …...(1)"


Put y = 0 in the equation of PQ for the coordinates of N.

"N \equiv (c(t + T),0)"

Similarly MN ="\sqrt{{(\dfrac{c(t + T)}2 – c(t + T)})^2 + {{(\dfrac{c(t + T)}{2tT} – 0}})^2}"

Simplifying we get MN "= [c(t + T)\sqrt{(1 + T^2t^2) / 2}] …...(2)"

From (1) & (2),

OM = MN